### v1

@@ Appendix A Lemmas
If the theorem holds for $t=1$, we obtain that at $s=1$ (notice $\left\|t_{0}H_{t_{0}s}\right\|\leq Mt_{0}$),
$\|G_{t_{0}}(i,j)\|\leq\exp(-Mt_{0})\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{(Mt_{0})^{k}}{k!}.$
Next we consider the case $t=1$.
II. Dyson series [7]. By variation of constants, we have
@@
In A.3 we show that the representation has an equivalent definition:
@@
Here $S_{n}$ is the set of all permutations of $[n]$. We use $P\{X\}$ to distinguish different ways of definitions. But they are in fact equivalent.
@@
$G_{1}=\exp(-M)\sum_{n=0}^{\infty}\frac{1}{n!}\hat{X}_{n}[X](1).$
Consider its finite truncation
$G_{1}=\exp(-M)\sum_{k=0}^{n}\frac{1}{k!}\hat{X}_{k}[X](1)+\exp(-M)\sum_{k=n+1}^{\infty}\frac{1}{k!}\hat{X}_{k}[X](1).$
Notice by the alternative representation,
@@
Since $0\preceq H_{t_{i}}\preceq MI$, it holds that $\left\|MI-H_{t_{i}}\right\|_{\rm op}\leq M$, and thus
$\|\hat{X}_{n}[X](1)\|_{\rm op}\leq M^{n}n!\int_{[0,1]^{n}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}=M^{n}.$ (A.2)
Since $H_{t}$ has block bandwidth $B$, the block bandwidth of $\hat{X}_{k}[X]$ is at most $kB$. So that if $|i-j|>nB\geq 0$, $\hat{X}_{k}[X](i,j)=0$ if $k\leq n$, and thus
$\|G_{1}(i,j)\|=~{}$ $\left\|\exp(-M)\sum_{k=n+1}^{\infty}\frac{1}{k!}\hat{X}_{k}[X](1)(i,j)\right\|$
[…]
@@
Here the first equality follows by discussion on the position of $x_{n}$. The third equality follows by integrating the variable $t_{k}$ and notice the constraint $t_{k+1}\leq t_{k}\leq t_{k-1}$. The forth equality follows by relabeling the index. The last equality follows by induction hypothesis. By symmetry, the relation holds for any $x_{i}$:
$\nabla_{x_{i}}\left(P\{X\}(t)\right)=-\left(a_{n}\prod_{j\neq i}(x-x_{j})\right)[X](t)=(\nabla_{x_{i}}P)[X](t),$
@@
Appendix B Proofs for 3.2
By 3.1 and 2.7, denote $\delta=\frac{sr!\kappa^{r}}{m}$,
$\max_{j\in V}W_{1}(\pi_{j},\pi_{r,j})\leq\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log\pi-\nabla_{j}\log\pi_{r}\right\|_{L^{1}(\pi_{r})}.$
Denote $x:=C^{-1/2}\tilde{x}$. By the chain rule,
@@
So that by B.1 (see below),
$\max_{j}\int|\nabla_{j}\log\pi(x)-\nabla_{j}\log\pi_{r}(x)|\pi_{r}(x)\mathrm{d}x$
@@
So that one obtain (denote $C_{\pi}=m^{-3/2}(sr!\kappa^{r})^{2}$)
$\max_{j}W_{1}(\pi_{j},\pi_{r,j})\leq C_{\pi}\max_{k}\int|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})|\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}.$ (B.1)
@@
|~$|\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}=|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})|^{2}+|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2},$
where we denote ~$\tilde{\nabla}_{r,k}=\nabla_{\tilde{x}_{r,k}},\tilde{\nabla}_{\bot,k}=\nabla_{\tilde{x}_{\bot,k}}$ and use ~$\tilde{\nabla}_{\bot,k}\log\tilde{l}_{r}(\tilde{x}_{r})=0$ as $\tilde{l}_{r}(\tilde{x}_{r})$ does not depend on $\tilde{x}_{\bot}$. Therefore,
|~$\begin{split}\int|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=\int|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}&\\ +\int|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=:~{}&\text{I}+\text{II}.\end{split}$ (B.2)
@@
$\begin{split}\text{I}=~{}&\int\text{\rm tr}\left[\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x}))^{\rm T}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\text{\rm tr}\left[\int\tilde{P}_{\bot,k}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x}))^{\rm T}\tilde{P}_{\bot,k}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right]\\ =~{}&\text{\rm tr}\left[\tilde{P}_{\bot,k}\widetilde{\mathcal{G}}_{k}\tilde{P}_{\bot,k}\right]=\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k}).\end{split}$ (B.3)
@@
$\text{II}=\int\left(\int|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}.$
@@
So that ~$\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})$ is the average of ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})$ with respect to $\tilde{\pi}_{0}(\tilde{x}_{\bot})$. Therefore, we can apply the Poincaré inequality for the function ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})$ with reference measure $\tilde{\pi}_{0}(\tilde{x}_{\bot})$ (which is a standard Gaussian) to obtain
|~$\int|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}$
@@
$\begin{split}\text{II}\leq~{}&\int\left(\int\sum_{j\in V}\|\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\|_{\rm F}^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}\\ \leq~{}&\int\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\tilde{P}_{\bot,j}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\widetilde{\mathcal{H}}_{j,k}\tilde{P}_{\bot,j}\right]=\sum_{v\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j}).\end{split}$ (B.4)
@@
$\max_{j}W_{1}(\pi_{j},\pi_{r,j})\leq~{}$ $C_{\pi}\max_{k}\left(\int|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right)^{1/2}$
@@
where we use Cauchy’s inequality in the first step.
@@ B.1 Lemma
Under 3.1, it holds that (denote $\kappa=\frac{M}{m}$)
$\max_{j\in V}\sum_{k\in V}\|C^{1/2}(j,k)\|\leq m^{-1/2}sr!\kappa^{r}.$ (B.5)
@@
For $j\neq k$, let $n=d_{G}(j,k)-1$, plug in $x=C^{-1}$ and notice $mI\preceq C^{-1}\preceq MI$, we obtain
@@
where $P_{n}\in\mathbb{P}_{n}$. By assumption, $C^{-1}(u,v)=0$ if $v\notin N_{u}$. Since $d_{G}(j,k)=n+1$, it holds
@@
as all the path connecting $j$ and $k$ has length no less than $d_{G}(j,k)>n$. Therefore,
$\|C^{1/2}(j,k)\|=~{}$ $M^{-1/2}\left\|\sum_{q=n+1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(I-\frac{C^{-1}}{M}\right)^{k}(j,k)\right\|$
@@
Here the second inequality follows the sparsity assumption, the third inequality follows A.4, and the last inequality follows from $\Gamma\left(r+\frac{1}{2}\right)\leq r!$ and $sr!\kappa^{r-1}\geq 1$. This completes the proof. ∎
Appendix C Proof of 4.1
By 2.7, $p_{\theta}\in\mathscr{P}$ is $\delta$-localized with $\delta=\frac{sr!\kappa^{r}}{m}$, and
$\begin{split}\max_{v\in V}W_{1}(p_{j},\hat{p}_{j})\leq~{}&\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log p-\nabla_{j}\log\hat{p}\right\|_{L^{1}(p)}\\ \leq~{}&\delta\cdot\left(\max_{j\in V}\mathbb{E}_{X\sim p}|s(X)-s_{\hat{\theta}_{*},v}(X)|^{2}\right)^{1/2}=\delta\cdot J_{\rm loc}^{1/2}(\hat{\theta}_{*}).\end{split}$ (C.1)
@@
$J_{\rm loc}(\hat{\theta}_{*})=~{}$ $\left(J_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\hat{\theta}_{*})\right)+\left(\hat{J}_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\theta_{*})\right)+\left(\hat{J}_{\rm loc}(\theta_{*})-J_{\rm loc}(\theta_{*})\right)$
[…]
@@
$\mathbb{P}_{N}f=\frac{1}{N}\sum_{i=1}^{N}f(X_{i}),\quad f_{\theta,j}(x)=2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(x)\right)+\left|s_{\theta,j}(x)\right|^{2}.$
Then $\hat{J}_{j}(\theta)-J_{j}(\theta)=(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta,j}$. For any $\theta,\phi\in\Theta$, it holds that
$|f_{\theta,j}-f_{\phi,j}$ $|\leq 2\left|\Delta_{jj}\log p_{\theta}-\Delta_{jj}\log p_{\phi}\right|+\left|\left|\nabla_{j}\log p_{\theta}\right|^{2}-\left|\nabla_{j}\log p_{\phi}\right|^{2}\right|$
@@
where we use $|N_{j}|\leq s+1$ and $\left|\nabla_{j}u_{\theta,k}(x)\right|\leq R$.
@@
By definition, $\sup_{u_{\theta}\in\mathcal{U}_{\theta}}\left\|u_{\theta}\right\|_{\mathcal{U}}\leq R$. Consider a $\varepsilon$-net $\mathcal{U}_{\varepsilon}=\{u_{i}\}_{n=1}^{N_{\varepsilon}}\subset\mathcal{U}_{\theta}$ of $\mathcal{U}_{\theta}$ with respect to $\left\|\cdot\right\|_{\mathcal{U}}$, i.e. $\forall u_{\theta}\in\mathcal{U}_{\theta}$, there exists some $u_{i}\in\mathcal{U}_{\varepsilon}$ s.t. $\left\|u_{\theta}-u_{i}\right\|_{\mathcal{U}}\leq\varepsilon$. Then
$\left\|f_{\theta,j}-f_{\theta_{i},j}\right\|_{L^{\infty}}\leq 2(s+1)R\sum_{k\in N_{j}}\left\|u_{\theta,k}-u_{\theta_{i},k}\right\|_{C^{2}}\leq 2(s+1)^{2}R\varepsilon.$
Here we take $\theta_{i}$ to be the parameter corresponding to $u_{i}$. Therefore,

### v2

@@ Appendix A Proof of Lemmas
If the theorem holds for $t=1$, then we obtain that at $s=1$ (note $\left\|t_{0}H_{t_{0}s}\right\|\leq Mt_{0}$),
$\|G_{t_{0}}(i,j)\|\leq\exp(-Mt_{0})\sum_{k=\mathsf{d}_{G}(i,j)}^{\infty}\frac{M^{k}t_{0}^{k}}{k!}.$
II. Dyson series solution. By variation of constants, we have
@@
In A.1 we show that the representation has an equivalent definition:
@@
Here $S_{n}$ is the set of all permutations of $[n]$. We use $P\{X\}$ to distinguish different ways of definitions. But they are in fact equivalent. Note also that the representation can be extended from polynomials to any analytic functions.
@@
$G_{1}=\exp(-M)\sum_{n=0}^{\infty}\frac{1}{n!}\hat{X}_{n}[X](1),$
@@
Here we use the alternative representation. We can truncate the Dyson series of $G_{1}$ as
Consider the off-diagonal entry $G_{1}(i,j)$. Take $n=\mathsf{d}_{G}(i,j)-1$, then since all the path in $G$ connecting $i$ and $j$ has length no less than $\mathsf{d}_{G}(i,j)>n$, it must hold that
$\forall 1\leq q\leq n,\quad\left[(MI-H_{t_{1}})\cdots(MI-H_{t_{q}})\right](i,j)=0~{}\Rightarrow~{}\hat{X}_{q}[X](1)(i,j)=0,$
Therefore,
$G_{1}(i,j)=\exp(-M)\sum_{q=n+1}^{\infty}\frac{1}{q!}\hat{X}_{q}[X](1)(i,j).$
Since $0\preceq H_{t_{k}}\preceq MI$, it holds that $\left\|MI-H_{t_{k}}\right\|_{\rm op}\leq M$, and thus
$\|\hat{X}_{n}[X](1)\|_{\rm op}\leq M^{n}n!\int_{[0,1]^{n}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}=M^{n}.$
$\Rightarrow~{}\|G_{1}(i,j)\|\leq\exp(-M)\sum_{q=n+1}^{\infty}\frac{M^{q}}{q!}=\exp(-M)\sum_{q=\mathsf{d}_{G}(i,j)}^{\infty}\frac{M^{q}}{q!}.$
@@
Here the first equality follows discussion on the position of $x_{n}$. The third equality follows integrating the variable $t_{k}$ and notice the constraint $t_{k+1}\leq t_{k}\leq t_{k-1}$. The forth equality follows relabeling the index and taking $t_{n}=0$. The last equality follows the induction hypothesis. By symmetry, the relation holds for any $x_{i}$:
$\nabla_{x_{i}}\left(P\{X\}(t)\right)=-\Big{(}a_{n}\prod_{j\neq i}(x-x_{j})\Big{)}[X](t)=(\nabla_{x_{i}}P)[X](t),$
@@
A.2 Lemma for 2.2
Lemma A.3.
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}\leq-MG_{t},$
where $\leq$ is in the entrywise sense, and $M\in\mathbb{R}^{b\times b}$ is $c$-diagonal dominant with non-positive off-diagonal entries, i.e. $\forall i,j\in[b],i\neq j$,
$\sum_{j\neq i}|M_{ij}|+c\leq M_{ii};\quad M_{ij}\leq 0.$
Then for any $t\geq 0$, it holds
$\|G_{t}\|_{\infty}\leq\mathrm{e}^{-ct}\|G_{0}\|_{\infty}.$
Proof.
Denote $G_{t}^{c}:=\mathrm{e}^{ct}G_{t}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{c}=\mathrm{e}^{ct}\left(\frac{\mathrm{d}}{\mathrm{d}t}G_{t}+cG_{t}\right)\leq\mathrm{e}^{ct}\left(-MG_{t}+cG_{t}\right)=\left(-M+cI\right)G_{t}^{c}.$
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{c}\mathbf{1}\leq\left(-M+cI\right)G_{t}^{c}\mathbf{1}.$
$\forall t\geq 0,\quad\|G_{t}^{c}\mathbf{1}\|_{\infty}\leq\|G^{c}_{0}\mathbf{1}\|_{\infty},$ (A.1)
which is a reformulation of the conclusion $\left\|G_{t}\right\|_{\infty}\leq\left\|G_{0}\right\|_{\infty}$ by noticing $G_{t}^{c}\in\mathbb{R}_{+}^{b}$ and
$\mathrm{e}^{ct}\|G_{t}\|_{\infty}=\|G_{t}^{c}\|_{\infty}=\max_{i}\sum_{j}G_{t}^{c}(i,j)=\|G_{t}^{c}\mathbf{1}\|_{\infty}.$
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{c}_{s}\mathbf{1})_{i}>0;\quad(G^{c}_{s}\mathbf{1})_{i}=\|G^{c}_{s}\mathbf{1}\|_{\infty}.$
On the other hand, notice by assumption on $M$,
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{c}_{s}\mathbf{1})_{i}\leq~{}$ $\left(-M_{ii}+c\right)(G^{c}_{s}\mathbf{1})_{i}+\sum_{j\neq i}(-M_{ij})(G^{c}_{s}\mathbf{1})_{j}$
[…]
Contradiction. This proves our result. ∎
Appendix B Proof of 4.1
Proof.
By 2.1, $\pi$ is $\delta$-localized with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$. 3.1 then implies
$\max_{i\in V}\mathsf{W}_{1}(\pi_{i},\pi_{r,i})\leq\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log\pi-\nabla_{j}\log\pi_{r}\right\|_{L^{1}(\pi_{r})}.$
Denote $x:=C^{-1/2}\tilde{x}$, $\nabla_{j}=\nabla_{x_{j}}$ and ~$\tilde{\nabla}_{j}=\nabla_{\tilde{x}_{j}}$. By the chain rule,
@@
By B.1 (see below), $\forall j,\sum_{k\in V}\|C^{1/2}(j,k)\|\leq m^{-1/2}S\nu!\kappa^{\nu}$. So that
$\max_{j}\int\|\nabla_{j}\log\pi(x)-\nabla_{j}\log\pi_{r}(x)\|\pi_{r}(x)\mathrm{d}x$
@@
Denote $C_{\pi}=m^{-3/2}(S\nu!\kappa^{\nu})^{2}$, then
$\max_{j}\mathsf{W}_{1}(\pi_{j},\pi_{r,j})\leq C_{\pi}\max_{k}\int\|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})\|\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}.$ (B.1)
@@
~$\|\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}=\|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})\|^{2}+\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2},$
where we denote ~$\tilde{\nabla}_{r,k}=\nabla_{\tilde{x}_{r,k}},\tilde{\nabla}_{\bot,k}=\nabla_{\tilde{x}_{\bot,k}}$ and use ~$\tilde{\nabla}_{\bot,k}\log\tilde{l}_{r}(\tilde{x}_{r})=0$ since $\tilde{l}_{r}(\tilde{x}_{r})$ does not depend on $\tilde{x}_{\bot}$. Therefore,
~$\begin{split}\int\|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=\int\|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}&\\ +\int\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=:~{}&\mathsf{I}+\mathsf{II}.\end{split}$ (B.2)
@@
$\begin{split}\mathsf{I}=~{}&\int\text{\rm tr}\left[\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x}))^{\rm T}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\text{\rm tr}\left[\int\tilde{P}_{\bot,k}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x}))^{\rm T}\tilde{P}_{\bot,k}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right]\\ =~{}&\text{\rm tr}\left[\tilde{P}_{\bot,k}\widetilde{\mathcal{G}}_{k}\tilde{P}_{\bot,k}\right]=\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k}).\end{split}$ (B.3)
@@
$\mathsf{II}=\int\left(\int\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}.$
@@
This implies that ~$\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})$ is the average of ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})$ with respect to $\tilde{\pi}_{0}(\tilde{x}_{\bot})$. Therefore, we can apply the Poincaré inequality for the function ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})$ with reference measure $\tilde{\pi}_{0}(\tilde{x}_{\bot})$ (which is a standard Gaussian) to obtain
~$\int\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}$
@@
$\begin{split}\mathsf{II}\leq~{}&\int\left(\int\sum_{j\in V}\|\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\|_{\rm F}^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}\\ \leq~{}&\int\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\tilde{P}_{\bot,j}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\widetilde{\mathcal{H}}_{j,k}\tilde{P}_{\bot,j}\right]=\sum_{j\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j}).\end{split}$ (B.4)
@@
$\max_{j}\mathsf{W}_{1}(\pi_{j},\pi_{r,j})\leq~{}$ $C_{\pi}\max_{k}\left(\int\|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right)^{1/2}$
@@
where we use Cauchy’s inequality in the first step. This completes the proof. ∎
@@ B.1 Lemma
Under 4.1, it holds that (denote $\kappa=\frac{M}{m}$)
$\max_{j\in V}\sum_{k\in V}\|C^{1/2}(j,k)\|\leq m^{-1/2}S\nu!\kappa^{\nu}.$ (B.5)
@@
For $j\neq k$, let $n=\mathsf{d}_{G}(j,k)-1$, plug in $x=C^{-1}$ and notice $mI\preceq C^{-1}\preceq MI$, we obtain
@@
where $P_{n}\in\mathbb{P}_{n}$. By assumption, $C^{-1}(i,j)=0$ if $j\notin\mathcal{N}_{i}$. Since $\mathsf{d}_{G}(j,k)=n+1$, it holds
@@
as all the path connecting $j$ and $k$ has length no less than $\mathsf{d}_{G}(j,k)>n$. Therefore,
$\|C^{1/2}(j,k)\|=~{}$ $M^{-1/2}\left\|\sum_{q=n+1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(I-\frac{C^{-1}}{M}\right)^{q}(j,k)\right\|$
@@
Here the second inequality follows the sparsity assumption, the third inequality follows A.2, and the last inequality follows $\Gamma\left(\nu+\frac{1}{2}\right)\leq\nu!$ and $s\nu!\kappa^{\nu-1}\geq 1$. This completes the proof. ∎
Appendix C Proof of 5.1
Proof.
By 2.1, $\pi$ is $\delta$-localized with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$. 3.1 then implies
$\begin{split}\max_{i\in V}\mathsf{W}_{1}(p_{i},\hat{p}_{i})\leq~{}&\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log p-\nabla_{j}\log\hat{p}\right\|_{L^{1}(p)}\\ \leq~{}&\delta\cdot\left(\max_{j\in V}\mathbb{E}_{X\sim p}\|s_{j}(X)-s_{\hat{\theta}_{*},j}(X)\|^{2}\right)^{1/2}=\delta\cdot J_{\rm loc}^{1/2}(\hat{\theta}_{*}).\end{split}$ (C.1)
@@
$\begin{split}J_{\rm loc}(\hat{\theta}_{*})=~{}&\left(J_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\hat{\theta}_{*})\right)+\left(\hat{J}_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\theta_{*})\right)+\left(\hat{J}_{\rm loc}(\theta_{*})-J_{\rm loc}(\theta_{*})\right)\\ \leq~{}&\left(J_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\hat{\theta}_{*})\right)+0+\left(\hat{J}_{\rm loc}(\theta_{*})-J_{\rm loc}(\theta_{*})\right)\\ \leq~{}&2\max_{\theta\in\Theta}\|\hat{J}_{\rm loc}(\theta)-J_{\rm loc}(\theta)\|\\ =~{}&2\max_{\theta\in\Theta}\left|\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(\hat{J}_{j}(\theta)+\lambda_{j}\right)-\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(J_{j}(\theta)+\lambda_{j}\right)\right|\\ \leq~{}&2\max_{\theta\in\Theta}\max_{j\in V}|\hat{J}_{j}(\theta)-J_{j}(\theta)|.\end{split}$ (C.2)
@@
$\mathbb{P}_{N}f=\frac{1}{N}\sum_{i=1}^{N}f(X_{i}),\quad f_{\theta,j}(x)=2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(x)\right)+\left\|s_{\theta,j}(x)\right\|^{2}.$
Then
$\hat{J}_{j}(\theta)-J_{j}(\theta)=(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta,j}.$
For any $\theta,\phi\in\Theta$, it holds that
$|f_{\theta,j}-f_{\phi,j}$ $|\leq 2|\Delta_{jj}\log p_{\theta}-\Delta_{jj}\log p_{\phi}|+|\left\|\nabla_{j}\log p_{\theta}\right\|^{2}-\left\|\nabla_{j}\log p_{\phi}\right\|^{2}|$
@@
where we use $|\mathcal{N}_{j}|\leq S+1$ and $\left\|\nabla_{j}u_{\theta,k}(x)\right\|\leq R$.
@@
By definition, $\sup_{u_{\theta}\in\mathcal{U}_{\theta}}\left\|u_{\theta}\right\|_{\mathcal{U}}\leq R$. Consider a $\varepsilon$-net $\mathcal{U}_{\varepsilon}=\{u_{\theta_{i}}\}_{n=1}^{N_{\varepsilon}}\subset\mathcal{U}_{\theta}$ of $\mathcal{U}_{\theta}$ with respect to $\left\|\cdot\right\|_{\mathcal{U}}$, i.e. $\forall u_{\theta}\in\mathcal{U}_{\theta}$, there exists some $u_{\theta_{i}}\in\mathcal{U}_{\varepsilon}$ s.t. $\left\|u_{\theta}-u_{\theta_{i}}\right\|_{\mathcal{U}}\leq\varepsilon$. Then
$\left\|f_{\theta,j}-f_{\theta_{i},j}\right\|_{L^{\infty}}\leq 2(S+1)R\sum_{k\in\mathcal{N}_{j}}\left\|u_{\theta,k}-u_{\theta_{i},k}\right\|_{C^{2}}\leq 2(S+1)^{2}R\varepsilon.$
Therefore,